OpenStudy (anonymous):

Circle S has an equation of (x + 16^)2 + (y – 9)^2 = 4. What is the center and radius of circle S? Center: (16, -9); Radius: 4 Center: (16, -9); Radius: 2 Center: (-16, 9); Radius: 4 Center: (-16, 9); Radius: 2

OpenStudy (helder_edwin):

the equation of a circle is \[ \large (x-h)^2+(y-k)^2=r^2 \] where the center is C(h,k) and radius is r.

OpenStudy (helder_edwin):

in your case u have \[ \large (x+16)^2+(y-9)^2=4 \] which can be rewritten as \[ \large (x-(-16))^2+(y-9)^2=2^2 \]

OpenStudy (anonymous):

sorry, i'm still a little lost. would you mind narrowing down the possible answers for me?

OpenStudy (helder_edwin):

i just did

jimthompson5910 (jim_thompson5910):

Compare \[ \large (x-h)^2+(y-k)^2=r^2 \] and \[ \large (x-(-16))^2+(y-9)^2=2^2 \] what is h, k and r?

OpenStudy (helder_edwin):

compare my first post and the last equation i wrote.

OpenStudy (helder_edwin):

@jim_thompson5910 put them together for u.